two minor league baseball player got a total of 398 hits. washington had 4 more hits than sanchez.whats the number of hits for each player.
1
Expert's answer
2012-09-27T12:18:30-0400
Let W be the number of washington's hits and S be the number of sanchez hits. Then we've got the following system:

W + S = 398;
S + 4 = W,

therefore

(S + 4) + S = 398 ==> 197,

and

W = 197 + 4 = 201.

So, washington has 201 hits and sanchez has 197 ones.